### v1

@@ 1. Restriction 1
    Block proposals can only extend a fastable block. Similarly, notarization shares, fast shares, and finalization shares are only sent for blocks that extend a fastable block.
@@ 2. Restriction 2
    Replicas move to the next round once a fastable block has been notarized.
@@ 3. Addition 1
    When replicas move to the next round, a fastable proof of the notarized block is broadcast.
@@ 4. Addition 2
    When a block is proposed, a fastable proof of the parent block is broadcast alongside the proposed block.
@@ 5. Addition 3
    Alongside the first round-$k$ notarization share, a fast share is broadcast.
@@ 6. Addition 4
    If a block has received $n-p$ fast shares, it is considered (FP-)finalized. Fast shares are combined into a fast finalization and broadcast.
Algorithms 1 and 2 contain the full FICC pseudocode.
@@
Fastable proofs (Addition 1) can be implemented naively by aggregating the fast shares for each block (Boneh et al., 2018). In periods of extreme asynchrony, no aggregation might be possible, from which follows a worst-case proof size of $12f+2p+$ 1 shares. In practice, however, we expect only a few different blocks to receive fast shares. The application of more targeted cryptographic primitives to achieve small fastable proofs in the worst case is left to future work.
@@
2:     Fast Internet Computer Consensus, instance ficc
3:     Code for round-$k$ at replica $u$
4:
5:Uses:
6:     RandomLeaderElectionOracle, instance oracle
7:     BestEffortBraodcast, instance broadcast
8:
9:Parameters:
10:     $\Delta_{prop}$triangleright Proposal Delay
11:     $\Delta_{notary}$triangleright Notarization Delay
12:     $k$triangleright Current round number
13:     $kMax$triangleright Highest round finalized so far
14:     $r_{u}$triangleright Rank assigned by the oracle
15:     $payload$triangleright Set of transactions for this round
16:     $b_{p}$triangleright Notarized and fastable round-($k-1$) block
17:
@@
19:     $fastSent:=\text{\tt False}$; $\triangleright$ Fast share sent
20:     $proposed:=\text{\tt False}$; $\triangleright$ Proposed a block
21:     $t_{0}:=clock()$; $\triangleright$ Time at the start of the round
22:     $D:=\emptyset$; $\triangleright$ Set of disqualified ranks
23:     $N:=\emptyset$; $\triangleright$ Set of blocks for which a notarization share was sent
24:
25:upon $\neg proposed\;\textbf{and}\;clock()\geq t_{0}+\Delta_{prop}(r_{u})$ do
26:     create a new round-k block:
27:     $b:=(k,u,hash(b_{p}),payload,\sigma_{u})$
28:     $proposed\leftarrow\text{\tt True}$
29:     broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s fastable proof
30:
31:upon exists a valid round-$k$ block $b$ of rank $r$ such that $b\notin N,\;\textbf{and}\;r\notin D\;\textbf{and}\;clock\geq t_{0}+\Delta_{notary}(r)\;\textbf{and}\;\nexists\textit{ valid round-$k$ block $b^{\prime}$ of rank $r^{\prime}$}\in[r]\setminus D$ do
32:     if $r_{u}\neq r$ then
33:         broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s fastable proof
34:     end if
35:     if some block in $N$ has rank $r$ then:
36:         $D\leftarrow D\cup\{r\}$;
37:     else
38:         $N\leftarrow N\cup\{b\}$
39:         if $\neg fastSent$ then:
40:              broadcast fast share for $b$;
41:              $fastSent\leftarrow\text{\tt True}$;
42:         end if
43:         broadcast notarization share for $b$
44:     end if
45:
Algorithm 1 FICC (Part 1)
@@
2:     combine notarization shares into a notarization for $b$
@@
5:     combine fast shares into a fastable proof
6:     broadcast notarization and fastable proof for $b$
@@
8:         broadcast finalization share for $b$
@@
13:     combine (fast) finalization shares into a (fast)
@@
19:     return $b$ extends a notarized and fastable round-$(k-1)$ block $b_{p}$ and $b$ is signed correctly
@@
Algorithm 2 FICC (Part 2)
@@ 8.1. Deadlock Freeness
The ICC protocol guarantees deadlock freeness as each honest replica receives at least one notarized block in each round. Therefore, each honest replica can start the following round and the block-tree keeps growing. However, in FICC  a block that was notarized is not guaranteed to be extended as, due to Restriction 1, in order for this to happen, the block must also be fastable.
The following theorem proves that regardless, deadlock freeness is not impacted. Intuitively, Restriction 2 makes sure that replicas only move to a higher round once a block can be extended, and Addition 1 guarantees that each replica does so eventually.
Theorem 8.1 ().
FICC satisfies deadlock freeness.
@@
We prove that the block-tree of each honest replica keeps growing in height by induction on the block-tree height $k$. Specifically, we show that for each round $k$, at least one notarized and fastable block will be added to each honest replica’s block-tree.
For the base case, note that all honest replicas agree on the root of the block-tree, the genesis block, by definition. This genesis block is notarized and finalized by definition too, which in turn makes it fastable by definition 7.4. Thus, in round 0, a notarized and fastable block is added to each node’s block-tree.
Now let’s assume that a notarized and fastable block exists in round $k$. We prove that each replica observes a notarized and fastable block in round $k+1$. First, note that all honest replicas will enter round $k+1$, by executing lines 49 to 55, as by the induction hypothesis they all have a block that can be extended.
Each honest replica will receive at least one round-(k+1) block, together with a notarization and fastable proof for the extended parent block (Addition 2). Thus, each honest replica will send a fast share for one block (Addition 3). Eventually, each replica is thus guaranteed to observe $n-f$ fast shares. Note that at the latest at this point, at least one block will be fastable, since either (i) one block receives more than $f+p$ fast shares, or (ii) other blocks receive more than $n-f-(f+p)>f+p$ fast shares. As no honest replica moves to a higher round without a fastable block being notarized (Restriction 2), and replicas keep notarizing blocks, eventually, one fastable block will be notarized. Honest nodes moving from one round to the next broadcast all the information needed for other honest nodes also to receive a notarized and fastable block (Addition 1).
@@
Any scenario in which notarized blocks for round-k exist before any of them are fastable can only occur due to message reordering. Hence Restriction 1 and Restriction 2 cannot cause latency concessions in the slow path if the communication channel precludes message reordering.
@@ 8.2. Safety
We guarantee that all honest replicas agree on the chain of finalized blocks. Suppose an honest replica finalizes block $b$ at height $k$ and let $pt$ be the path in the block-tree from the genesis to $b$. Suppose another honest replica finalizes block $b^{\prime}$ at height $k^{\prime}$, where $k^{\prime}\geq k$, and let $pt^{\prime}$ be the path in the block-tree from genesis to $b^{\prime}$.
Requiring that all honest replicas agree on the finalized blocks is equivalent to requiring that $pt$ be a prefix of $pt^{\prime}$. This is guaranteed as long as, if a block is finalized (either IC-finalized or FP-finalized), any other block at the same height can be neither finalized nor extended. Thus, if a block at height $k$ is IC-finalized, there must not be another block at height $k$ which is notarized. Similarly, if a block at height $k$ is FP-finalized, there must not be another block at height $k$ which is fastable.
Lemma 8.2 ().
Let $n\geq 3f+2p+1$ , in any round, if a block $b$ is IC-finalized, then FICC guarantees that no block $b^{\prime}$, $b^{\prime}\neq b$ is notarized for any replica.
@@
As no changes have been made to FICC in terms of notarization, this statement follows directly from lemma 4.1. ∎
Therefore, any other block at height $k$ will be neither extended nor finalized (as both cases require it to be at least notarized). Similarly, let $b$ be a block at height $k$ which is FP-finalized, we prove that $b$ is the only fastable block at height $k$.
Lemma 8.3 ().
Let $n\geq 3f+2p+1$ , in any round, if a block $b$ is FP-finalized, then FICC guarantees that no block $b^{\prime}$, $b^{\prime}\neq b$ is fastable for any replica.
@@
Assume towards contradiction that there exists an FP-finalized block $b$ and a fastable block $b^{\prime}$ at height $k$. By the definition of FP-finalization a replica observed $|FS(b)|\geq n-p\geq 3f+p+1$ fast shares for $b$. Thus, all blocks different from $b$ have received at most $f+p$ fast shares in total. In particular $|FS(b^{\prime})|\leq f+p$. This means that condition (1) of definition 7.4 is not met and hence condition (2) must hold for $b^{\prime}$ to be fastable: $sum(k)-{\left|FS(m(k))\right|}>f+p$.
  • Let’s assume $m(k)=b$, then $sum(k)>f+p+|FS(b)|$, which implies that more than $p$ replicas have not sent a fast share for $b$, thus reaching a contradiction.
  • Instead, let’s assume $m(k)=b^{\prime}$, $b^{\prime}\neq b$. Then ${\left|FS(b)\right|}\leq{\left|FS(b^{\prime})\right|}$ by the definition of $m(k)$. Thus $sum(k)>f+p+|FS(b^{\prime})|$ implies $sum(k)>f+p+|FS(b)|$, which is a contradiction since there can be at most $p$ honest replicas not sending a fast share for $b$.
Figure 5. The rank-1 (r-1) block is fastable (Fa) for replica $u$, while the rank-0 (r-0) block is FP-finalized for replica $v$. Lemma 8.3 proves that this situation cannot arise.
Theorem 8.4 ().
FICC satisfies safety.
@@
We start by noting that FP-finalization implies notarization, hence IC- and FP-finalization cannot happen on different blocks of the same round. Lemma 8.2 proves that, if a block is IC-finalized, then there is no other notarized block at the same height. Likewise, lemma 8.3 proves that, if a block is FP-finalized, then there is no other fastable block at the same height.
In both cases, since only notarized and fastable blocks are extended, if a block at height $k$ is finalized, no other blocks at height $k$ will be extended and no fast, notarization, or finalization shares will be sent for these blocks (Restriction 1). This guarantees that whenever a block might get finalized all replicas only extend this block. Thus finalized blocks are always part of the same blockchain. ∎
@@ 8.3. Liveness
Theorem 8.5 ().
FICC satisfies liveness.
@@
Let $T$ be the time when the first honest replica $u$ enters round $k$. Suppose that the rank-0 replica $q$ of round $k$ is honest. Moreover, suppose that the network is $\delta$-synchronous between time $T$ and $T+2\delta+\Delta_{prop}(0)$. Furthermore, assume $\delta\leq\Delta_{bnd}$, which implies that $\Delta_{notary}(1)\geq 2\delta+\Delta_{prop}(0)$, with the functions $\Delta_{notary}$ and $\Delta_{prop}$ as defined in section 4.
We want to prove that under these assumptions, eventually, each honest replica will see replica $q$’s proposed block as finalized.
Before the honest replica $u$ enters round $k$ at time $T$, the replica $u$ notarized a fastable block in rounds $11,\dots,k-$ 1 and broadcast the notarization and fastable proof (Addition 1). By the synchrony assumption, the other replicas, $q$ included, receive the notarization and fastable proof for rounds $11,\dots,k-$ 1 by time $T+\delta$, and enter round $k$. Thus, replica $q$ broadcasts its rank-0 proposal by time $T+\delta+\Delta_{prop}(0)$, and other replicas will receive it by $T+2\delta+\Delta_{prop}(0)$. By assumption no other honest replica enters round $k$ before time $T$ and as $\Delta_{notary}(1)\geq 2\delta+\Delta_{prop}(0)$, every honest replica will broadcast a fast share and a notarization share for the rank-0 block (Addition 3). Therefore, all honest replicas eventually receive at least $n-f$ fast shares for the same rank-0 block. When $p=f$, these fast shares are enough to FP-finalize the block proposed by replica $q$ (Addition 4). When $p<f$, ${\lfloor\frac{n+f}{2}\rfloor}+1\leq n-f$ notarization shares are sufficient to notarize the block instead. As argued previously, honest replicas will not have notarized another block and thus will broadcast a finalization share. Eventually, after an additional communication round honest replicas will receive ${\lfloor\frac{n+f}{2}\rfloor}+1\leq n-f$ finalization shares, which leads to IC-finalization. ∎

### v2

@@ 1. Restriction 1
    Block proposals can only extend an unlocked block. Similarly, notarization votes, fast votes, and finalization votes are only sent for blocks that extend an unlocked block. (We change the validity condition on 18.)
@@ 2. Restriction 2
    Replicas move to the next round once an unlocked block has been notarized, and they have sent a fast vote (4).
@@ 3. Addition 1
    When replicas move to the next round, an unlock proof of the notarized block is broadcast (32).
@@ 4. Addition 2
    When a block is proposed, an unlock proof of the parent block and a fast vote for the current block is broadcast alongside the proposed block (32 and 29).
@@ 5. Addition 3
    When the first round $k$ notarization vote is broadcast, a fast vote for the same block is broadcast alongside it (43).
@@ 6. Addition 4
    If a rank-0 block has received $n-p$ fast votes, it is considered FP-finalized (12). Fast votes are aggregated into a fast finalization and broadcast (13 and 14).
Definition 7.6 (Unlock Proof).
An unlock proof (Addition 1) is the collection of valid fast votes, that prove that $b$ is unlocked according to Definition 7.5.
Unlock proofs can be implemented naively by aggregating the fast votes for each block using BLS multi-signatures (Boneh et al., 2018). In the worst case, condition 2) of definition 7.5 might only be met after receiving $12f+2p+$ 1 fast votes, which might attest $f+p+2$ unique blocks, leaving little improvements from aggregation. In practice, however, we expect only a few different blocks to receive fast votes. The description of a more targeted mechanism to achieve small unlock proofs in the worst case is left to future work.
@@
It is possible to omit sending a corresponding notarization vote when a fast vote is sent. A notarization then consists of two multi-signatures, on for notarization and one for fast votes. For the sake of simplicity, we do not consider this version of Banyan in our description and analysis.
@@
2:     Pseudocode for round $k$ at replica $u$
3:
4:Uses:
5:     RandomBeacon, instance beacon
6:     BestEffortBraodcast, instance broadcast
7:
8:Parameters:
9:     $\Delta_{prop}$triangleright Proposal delay
10:     $\Delta_{notary}$triangleright Notarization delay
11:     $k$triangleright Current round number
12:     $kMax$triangleright Highest round finalized so far
13:     $payload$triangleright Set of transactions for this round
14:     $r_{u}$triangleright Rank, permutation derived from beacon
15:     $b_{p}$triangleright Notarized and unlocked round ($k-1$) block
16:
@@
18:     $fastVoteSent\leftarrow\text{\tt False}$;
19:     $proposed\leftarrow\text{\tt False}$; $\triangleright$ Proposed a block
20:     $t_{0}\leftarrow clock()$; $\triangleright$ Time at the start of the round
21:     $D\leftarrow\emptyset$; $\triangleright$ Set of disqualified ranks
22:     $N\leftarrow\emptyset$; $\triangleright$ Set of blocks for which a notarization vote was sent
23:
24:upon $\neg proposed\;\textbf{and}\;clock()\geq t_{0}+\Delta_{prop}(r_{u})$ do
25:     create a new round k block:
26:     $b\leftarrow(k,u,hash(b_{p}),payload,{signature}_{u})$
27:     $proposed\leftarrow\text{\tt True}$
28:     if $r_{u}=0$ then
29:         broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s unlock proof, fast vote for $b$
30:         $fastVoteSent\leftarrow\text{\tt True}$
31:     end if
32:     broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s unlock proof
33:
34:upon exists a valid round $k$ block $b$ of rank $r$ such that $b\notin N,\;\textbf{and}\;r\notin D\;\textbf{and}\;clock\geq t_{0}+\Delta_{notary}(r)\;\textbf{and}\;\nexists\textit{ valid round $k$ block $b^{\prime}$ of rank $r^{\prime}$}\in[r]\setminus D$ do
35:     if $r_{u}\neq r$ then
36:         broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s unlock proof
37:     end if
38:     if some block in $N$ has rank $r$ then:
39:         $D\leftarrow D\cup\{r\}$;
40:     else
41:         $N\leftarrow N\cup\{b\}$
42:         if $\neg fastVoteSent$ then:
43:              broadcast fast and notarization vote for $b$;
44:              $fastVoteSent\leftarrow\text{\tt True}$;
45:         else
46:              broadcast notarization vote for $b$
47:         end if
48:     end if
49:
Algorithm 1 Banyan (Part 1)
@@
2:     combine notarization votes into a notarization for $b$
@@
5:     combine fast votes into a unlock proof
6:     broadcast notarization and unlock proof for $b$
@@
8:         broadcast finalization vote for $b$
@@
13:     combine (fast) finalization votes into a (fast)
@@
19:     return $b$ extends a notarized and unlocked round $(k-1)$ block $b_{p}$ and $b$ is signed correctly and contains a fast vote from the proposer if $b.rank=0$
@@
Algorithm 2 Banyan (Part 2)
@@ 8. Protocol Analysis
In this section, we present proof sketches that explain the correctness of Banyan. These are not complete proofs but are intended to provide a conceptual understanding of the protocol’s security.
@@ 8.1. Deadlock Freeness
The ICC protocol guarantees deadlock freeness as each honest replica receives at least one notarized block in each round. Therefore, each honest replica can start the next round and the block-tree keeps growing. However, in Banyan  a block that was notarized is not guaranteed to be extended as, due to Restriction 1, in order for this to happen, the block must also be unlocked.
The following lemma shows how, regardless, deadlock freeness is not impacted. Intuitively, Restriction 2 makes sure that replicas only move to a higher round once a block can be extended, and Addition 1 guarantees that each replica does so eventually.
Lemma 8.1.
If each honest replica broadcasts a fast vote in round $k$, then Banyan guarantees that each honest replica eventually observes at least one unlocked block.
@@
Assume towards contradiction that in round $k$ no block is unlocked for an honest replica $u$. Consider the set $\mathcal{S}$, defined to be the support for leader blocks different from $\textsc{Max}(k)$, i.e.,
$\mathcal{S}={\textsc{supp}(\textsc{nonMaxBlocks}(k)\setminus\textsc{nonLeaderBlocks}(k))}$
If $\mathcal{S}=\emptyset$, a contradiction is reached immediately, as
${\left|\textsc{supp}(\textsc{max}(k))\right|}+{\left|\textsc{supp}(\textsc{nonLeaderBlocks}(k))\right|}$
$={\left|\textsc{supp}(\textsc{blocks}(k))\right|}>n-f>f+p$
and thus by Item 1 of Definition 7.5 $\textsc{max}(k)$ would be unlocked. (This corresponds to rounds with an honest leader.)
Instead, if $\mathcal{S}\neq\emptyset$, there are at least two rank-0 blocks. By Addition 2 (29), they each contain a fast vote from the (Byzantine) leader. Since $u$ also receives fast votes from $n-f\geq 2f+2p-1$ honest replicas, at least $12f+2p+$ 1 fast votes will be received, i.e., ${\left|\textsc{supp}(\textsc{max}(k)\cup\textsc{nonMaxBlocks}(k))\right|}=2f+2p+1$. By the pigeonhole principle, either ${\left|\textsc{supp}(\textsc{max}(k))\right|}\geq f+p+1$ or ${\left|\textsc{supp}(\textsc{nonMaxBlocks}(k))\right|}\geq f+p+1$. In both cases at least one block is unlocked according to Definition 7.5, thus leading to a contradiction. ∎
Theorem 8.2.
Banyan satisfies deadlock freeness.
Proof.
We prove that each round eventually terminates, and that the block-tree of each honest replica keeps growing in height by induction on the block-tree height $k$. Specifically, we show that for each round $k$, at least one notarized and unlocked block will be added to each honest replica’s block-tree.
Base case: All honest replicas agree on the root of the block-tree, the genesis block, which is also notarized and finalized by definition. By Definition 7.5 the genesis block is thus unlocked. Thus, in round 0, a notarized and unlocked block is added to each replica’s block-tree.
Induction step: Assuming a notarized and unlocked block exists in round $k$, we prove that each replica observes a notarized and unlocked block in round $k+1$. By the induction hypothesis, all replicas will broadcast a fast vote and thus execute the procedure that starts on 4. Thus, all honest replicas will enter round $k+1$.
Each honest replica will receive at least one round $k+1$ block, together with a notarization and unlock proof for the extended parent block (Addition 2, 32). Thus, each honest replica will send a fast vote for one block (Addition 3, 43). Thus, by Lemma 8.1, at least one block round $k+1$ block will be unlocked. As no honest replica moves to a higher round without an unlocked block being notarized (Restriction 2, 4), and replicas keep notarizing blocks, one unlocked block will be notarized. Finally, honest replicas moving from one round to the next broadcast the notarization and unlock proof (Addition 1, 6), guaranteeing that all honest replicas can add a notarized and unlocked block to their block-tree. ∎
@@
Banyan is at least as fast as ICC. Any scenario in which a notarized block for round k exists before it is unlocked can only occur due to message reordering. Hence, Restriction 1 and Restriction 2 cannot cause latency concessions if the communication channel precludes message reordering (which is the case in practice when TCP/QUIC is used).
@@ 8.2. Safety
ICC derives its safety from the fact that in rounds with an explicetly finalized block, no other blocks can be notarized. We start our way towards sketching the safety proof of Banyan by showing that an equivalent property holds for Banyan.
Lemma 8.3.
If a round $k$ block $b$ is SP-finalized, then Banyan guarantees that no round $k$ block $b^{\prime}$, $b^{\prime}\neq b$ is notarized for any replica.
@@
Let block $b$ be SP-finalized by some replica, and let block $b^{\prime}$, $b^{\prime}\neq b$ be notarized. A replica must have received a quorum of at least ${\lceil\frac{n+f+1}{2}\rceil}$ finalization votes for $b$, ${\lceil\frac{n-f+1}{2}\rceil}$ of which are from honest replicas. Moreover, some replica must have received a quorum of ${\lceil\frac{n+f+1}{2}\rceil}$ notarization votes for $b^{\prime}$, ${\lceil\frac{n-f+1}{2}\rceil}$ of which are from honest replicas. The two quorums must be disjoint, since each honest replica sends a finalization vote for $b$ only if it did not send any notarization vote for other blocks (in particular $b^{\prime}$) at height $k$ (7). Thus, there must be at least ${\lceil\frac{n-f+1}{2}\rceil}+{\lceil\frac{n-f+1}{2}\rceil}\geq n-f+1$ honest replicas. By definition, there are only $n-f$ honest replicas, hence one honest replica must occur in both quorums, a contradiction. ∎
Therefore, any other block at height $k$ will be neither extended nor finalized (as both cases require it to be at least notarized). Similarly, let $b$ be a block at height $k$ which is FP-finalized, we prove that $b$ is the only unlocked block at height $k$.
Lemma 8.4.
If a round $k$ block $b$ is FP-finalized, then Banyan guarantees that no round $k$ block $b^{\prime}$, $b^{\prime}\neq b$ is unlocked for any replica.
@@
Assume towards contradiction that there exists an FP-finalized block $b$ and an unlocked block $b^{\prime}$, $b^{\prime}\neq b$ at height $k$. For $b^{\prime}$ to be unlocked, either Condition 1 or Condition 2 of Definition 7.5 must be satisfied.
If Condition 1 is satisfied, then
${\left|\textsc{supp}(b^{\prime})\right|}+{\left|\textsc{supp}(\textsc{nonLeaderBlocks}(k))\right|}>f+p$
This implies that blocks different from $b$ have received more than $f+p$ fast votes, more than $p$ of which coming from honest replicas. By 42, 43 and 44, each honest replica broadcasts at most one fast vote per round. Thus $b$ received less than $n-p$ fast votes, a contradiction, as $n-p$ are required for FP-finalization (12).
If Condition 2 is satisfied, then
${\left|\textsc{supp}(\textsc{nonMaxBlocks}(k))\right|}>f+p$
It follows that ${\left|\textsc{supp}(\textsc{max}(k))\right|}<n-p$. This implies that no block can be FP-finalized, a contradiction. ∎
Theorem 8.5.
Banyan satisfies safety.
@@
Towards contradiction, assume an honest replica finalizes round $k$ block $b$, while another honest replica finalizes round $k$ block $b^{\prime}$, $b^{\prime}\neq b$. Without loss of generality, we assume $b$ was explicitly finalized.
Consider the case where $b$ and $b^{\prime}$ were both explicitly finalized. SP-and FP-finalization imply the reception of ${\lceil\frac{n+f+1}{2}\rceil}$ unique notarization and fast votes for a block, respectively. This implies the existence of two respective Byzantine quorums, a contradiction (Cachin et al., 2011). Thus, without loss of generality, assume $b^{\prime}$ was implicitly finalized. If $b$ was SP-finalized, by Lemma 8.3 $b^{\prime}$ cannot have been notarized, and thus no honest replica would have extended $b^{\prime}$. Instead, if $b$ was FP-finalized, by Lemma 8.4 then $b^{\prime}$ was not unlocked. Thus, honest replicas would have sent fast, notarization, or finalization votes for a block that was not unlocked, a contradiction (see 34 and the validity definition on 18). ∎
@@ 8.3. Liveness
Theorem 8.6.
Banyan satisfies liveness.
@@
Let $T$ be the time when the first honest replica $u$ enters round $k$. Suppose that the leader $l$ of round $k$ is honest and proposes block $b$. Moreover, suppose that the network is $\delta$-synchronous between time $T$ and $T+2\delta$. Furthermore, assume $\delta\leq\Delta$, which implies that $\Delta_{notary}(1)\geq 2\delta$, with the functions $\Delta_{notary}$ and $\Delta_{prop}$ as defined in Section 4. We want to prove that under these assumptions, eventually, each honest replica will finalize block $b$.
Before the honest replica $u$ enters round $k$ at time $T$, the replica $u$ notarized an unlocked block in rounds $11,\dots,k-$ 1 and broadcast the notarization and unlock proof (Addition 1, 6). By the synchrony assumption, the other replicas, $l$ included, receive the notarization and unlock proof for rounds $11,\dots,k-$ 1 by time $T+\delta$, and enter round $k$.
Note that $\Delta_{prop}(0)=0$. Thus, replica $l$ broadcasts $b$ by time $T+\delta$, and other replicas will receive it by $T+2\delta$. By assumption, no other honest replica enters round $k$ before time $T$ and as $\Delta_{notary}(1)\geq 2\delta$, every honest replica will broadcast a fast vote and a notarization vote for $b$ (Addition 3, 43). Therefore, all honest replicas eventually receive at least $n-f$ fast votes for $b$. When $p=f$, these fast votes are enough to FP-finalize $b$ (Addition 4, 12). When $p<f$, ${\lceil\frac{n+f+1}{2}\rceil}\leq n-f$ notarization votes are sufficient to notarize $b$ instead. As argued previously, honest replicas will not have notarized another block and thus will broadcast a finalization vote (8). Eventually, after an additional communication round, honest replicas will receive ${\lceil\frac{n+f+1}{2}\rceil}\leq n-f$ finalization votes, which leads to SP-finalization (12). ∎